OpenStudy (anonymous):

Choose the best definition for the following term: exponent. a.The number listed first when combining a variable and a number b.The number or value being raised to a power c.A letter that holds the place of some unknown value in mathematics d.Shorthand that tells how many times the base is used as a factor, or how many times the base is multiplied by itself

OpenStudy (debbieg):

What do you think? :)

OpenStudy (anonymous):

a.

OpenStudy (debbieg):

The number listed first when combining a variable and a number means something like the "3" in the expression: 3x That's not an exponent, that's called a "coefficient".

OpenStudy (debbieg):

Think about what an exponent MEANS. \(\Large 2^3=2\cdot 2\cdot 2\) \(\Large x^5=x\cdot x\cdot x \cdot x \cdot x\) What did I DO there? In the first one, I lined up three 2's, and multiplied them. In the second one, I lined up five x's, and multiplied them.

OpenStudy (debbieg):

By the way, the "2" and the "x" in each expression is called the "base".

OpenStudy (anonymous):

so it's d

OpenStudy (debbieg):

Yes ma'am. :)